# Constructing Concise Characteristic Samples for Acceptors of Omega Regular Languages

### v3

@@ 8.4. The learning algorithm $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IPA}}}$
Let $C$ be any node of $\mathcal{F}^{*}(\mathcal{P})$, and let $D$ be a child of $C$ in $\mathcal{F}^{*}(\mathcal{P})$. Then $D$ is an SCC, $D\subseteq C$, the parity of $D$ is opposite to the parity of $C$, and $D$ is maximal in the subset ordering with these properties, by property (3) of Thm. 8.5. In the sample $T_{Acc}^{\mbox{\scriptsize{IPA}}}$ there is an example $(u(v)^{\omega},l)$ with $D={\textsl{inf}}(u(v)^{\omega})$, so $D$ is an element of $Z$, and will be made a child of $C$ in $\mathcal{F}^{\prime}$ because $D\subseteq C$, the label $l$ is the opposite of the label of $C$, and $D$ is maximal in $Z$ with these properties. Conversely, if $D$ is made a child of $C$ in $\mathcal{F}^{\prime}$, then $D\subseteq C$, the label of $D$ is opposite to the label of $C$ (that is, they are of opposite $\kappa$-parity), and $D$ is maximal in $Z$ with these properties. This implies $D$ is a child of $C$ in $\mathcal{F}^{*}(\mathcal{P})$, by property (3) of Thm. 8.5.
By induction, $\mathcal{F}^{\prime}$ is equal to $\mathcal{F}^{*}(\mathcal{P})$, and therefore $\kappa^{\prime}$ is equal to the canonical coloring $\kappa^{*}$. Then the IPA $(\mathcal{M},\kappa^{\prime})$ is equivalent to $\mathcal{P}$, by Thm. 8.6. Because $T$ is consistent with $\mathcal{P}$, the IPA $(\mathcal{M},\kappa^{\prime})$ is returned by $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IPA}}}$. ∎
@@
Assume the sample $T$ is consistent with $\mathcal{P}$ and subsumes $T_{L}$. By Thm. 6.2, the automaton $\mathcal{M}$ is isomorphic to the automaton of $\mathcal{P}$. By Thm. 8.7, the acceptor $(\mathcal{M},\kappa^{\prime})$ is equivalent to $\mathcal{P}$. Because it is consistent with $T$, the IPA $(\mathcal{M},\kappa^{\prime})$ is the acceptor returned by $\mathbf{L}^{\mbox{\scriptsize{IPA}}}$. ∎
9. The sample $T_{Acc}$ and learning for IRA
@@ 9.2. An ordering on sets of states
Let $\mathcal{M}=\langle\Sigma,Q,q_{\iota},\delta\rangle$ be a deterministic complete automaton in which every state is reachable from the initial state $q_{\iota}$. Recall from subsection 6.2 that $\textit{access}(q)$ is the shortlex least string $s\in\Sigma^{*}$ such that $\delta(q_{\iota},s)=q$.
@@ 9.3. The learning algorithm $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IRA}}}$
Otherwise, $z_{i}$ causes the update to $\beta_{k}$, and among the pairs in $S_{k}$ that are added to $\beta_{k}$ is at least one that $z_{i}$ satisfies. To see this, note that $z_{i}$ must satisfy some pair in $\alpha$, say $(q,B_{j})$. Thus, $q\in{\textsl{inf}}(z_{i})$ and $B_{j}\cap{\textsl{inf}}(z_{i})=\emptyset$. The pair $(q,B)$ where $B=Q\setminus{\textsl{inf}}(z_{i})$ has $B_{j}\subseteq B$. Thus any $\omega$-word that satisfies $(q,B)$ will also satisfy $(q,B_{j})$. Because $T$ is consistent with $\mathcal{R}$, there can be no negative example in $T$ satisfying $(q,B)$, so pair $(q,B)$ is part of $S_{k}$ and is added to $\beta_{k}$. The word $z_{i}$ (and the word $w$) satisfies $(q,B)$ and is therefore accepted by $(\mathcal{M},\beta_{k+1})$ and every subsequent hypothesis, including $(\mathcal{M},\beta)$. ∎
@@ 9.5. Correctness of $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IRA}}}$
Now assume that the sample $T$ is consistent with $\mathcal{R}$ and subsumes $T_{L}$. Then by Thm. 6.2, the automaton $\mathcal{M}$ is isomorphic to the automaton of $\mathcal{R}$. By Thm. 9.4, the acceptor returned by $\mathbf{T}^{\mbox{\scriptsize{IRA}}}$ is an IRA $(\mathcal{M},\beta)$ is equivalent to $\mathcal{R}$, and this is the acceptor also returned by $\mathbf{L}^{\mbox{\scriptsize{IRA}}}$. ∎
@@ 10.1. Computing $T_{Aut}$
If we have a procedure to solve the inclusion problem, at most two calls to it will solve the equivalence problem. We describe polynomial time algorithms to solve the inclusion problem for DBAs, DCAs, and DPAs in Section 11 and 12, for DRAs and DSAs in Section 13, and for DMAs in Section 14. Referring to those sections, we obtain polynomial time algorithms to solve the equivalence problem for DBAs, DCAs, and DPAs from Thm. 11.1 and Thm. 12.2, for DRAs and DSAs from Thm. 13.2, and for DMAs from Thm. 14.6. Thus, we have the following.
@@
Given an acceptor $\mathcal{A}$ of type IBA, ICA, IPA, IRA, ISA, or IMA, the sample $T_{Aut}$ for the automaton portion of $\mathcal{A}$ may be computed in polynomial time.
@@ 10.2. Computing $T_{Acc}$
Given an IBA $\mathcal{B}=\langle\Sigma,Q,q_{\iota},\delta,F\rangle$ with automaton $\mathcal{M}$, the sample $T_{Acc}^{\mbox{\scriptsize{IBA}}}$ described in subsection 7.2 is computed as follows. For each SCC $C\in\textit{maxSCCs}(Q\setminus F)$, let $u(v)^{\omega}=\mathbf{Witness}(C,\mathcal{M})$ and include the negative example $(u(v)^{\omega},0)$ in $T_{Acc}^{\mbox{\scriptsize{IBA}}}$.
@@
Given an IPA $\mathcal{P}$ with automaton $\mathcal{M}$, the computation of $T_{Acc}^{\mbox{\scriptsize{IPA}}}$ proceeds as described in subsection 8.3. That is, the canonical forest $\mathcal{F}^{*}(\mathcal{P})$ is computed in polynomial time, and for each node $C$ in the forest, $u(v)^{\omega}=\mathbf{Witness}(C,\mathcal{M})$ is computed and $(u(v)^{\omega},l)$ is added to $T_{Acc}^{\mbox{\scriptsize{IPA}}}$, where $l$ is the label of node $C$ in the canonical forest. ∎
@@
Given an IRA $\mathcal{R}=(\mathcal{M},\alpha)$, the computation of $T_{Acc}$ proceeds as described in subsection 9.4. At each stage of the computation, it is necessary to find an $\omega$-word $u(v)^{\omega}$ with the $\preceq$-largest ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega}$) that is accepted by $(\mathcal{M},\alpha)$ and rejected by $(\mathcal{M},\gamma_{k})$. Thm. 13.2 gives a polynomial time algorithm that not only tests the inclusion of two DRAs, but returns a witness $u(v)^{\omega}$ with the $\preceq$-largest ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$ in the case of non-inclusion, because $\mathcal{M}$ is isomorphic to $\mathcal{M}\times\mathcal{M}$. ∎
@@
Given an IMA $\mathcal{A}=(\mathcal{M},{\mathcal{F}})$, the sample $T_{Acc}^{\mbox{\scriptsize{IMA}}}$ described in subsection 7.1 is computed as follows. For each $F\in{\mathcal{F}}$ determine whether $F$ is a reachable SCC of $\mathcal{M}$, and if so, compute $u(v)^{\omega}=\mathbf{Witness}(F,\mathcal{M})$ and add $(u(v)^{\omega},1)$ to the sample $T_{Acc}^{\mbox{\scriptsize{IMA}}}$. ∎
@@
Let $\mathcal{A}$ be an IBA, IPA, IRA, or IMA accepting the $\omega$-language $L$. Then the characteristic sample $T_{L}$ for $\mathcal{A}$ may be computed in polynomial time in the size of $\mathcal{A}$.
@@
By Thm. 10.1, the sample $T_{Aut}$ can be computed in polynomial time in the size of $\mathcal{A}$, and by Proposition 10.2, 10.3, 10.4, or 10.5 the sample $T_{Acc}^{\mbox{\scriptsize{IBA}}}$, $T_{Acc}^{\mbox{\scriptsize{IPA}}}$, $T_{Acc}^{\mbox{\scriptsize{IRA}}}$, or $T_{Acc}^{\mbox{\scriptsize{IMA}}}$ can also be computed in polynomial time in the size of $\mathcal{A}$. ∎
Note that Thm. 10.6 does not imply that the class $\mathbb{IXA}$ for $\mathbb{X}\in\{\mathbb{B},\mathbb{P},\mathbb{R},\mathbb{M}\}$ is efficiently teachable, since this class also has representations by non-isomorphic automata.
@@ 11. Inclusion algorithms
We show that there are polynomial time algorithms for the inclusion problem for DBAs, DCAs, DPAs, DRAs, DSAs, and DMAs. Recall that two calls to an inclusion algorithm suffice to solve the equivalence problem. By Claim 4, the inclusion and equivalence problems for DCAs are efficiently reducible to those for DBAs, and vice versa. Also, by Claim 1, the inclusion and equivalence problems for DBAs are efficiently reducible to those for DPAs. Thus it suffices to consider the inclusion problem for DPAs, DRAs, and DMAs.
@@
For the inclusion problem for DBAs, DCAs, and DPAs, Schewe [Sch10, Sch11] gives the following result.
@@
The inclusion problems for DBAs, DCAs, and DPAs are in NL.
Because NL (nondeterministic logarithmic space) is contained in polynomial time, this implies the existence of polynomial time inclusion and equivalence algorithms for DBAs, DCAs, and DPAs. For the sake of completeness, and to address the problem of returning a witness we include a proof sketch.
@@
For inclusion of DRAs and DMAs, [CDK93] provide a reduction to the problem of model checking a formula in the temporal logic CTL. While the complexity of model checking CTL formulas is in general PSPACE-hard, for the fair-CTL fragment, [EL87] provides a model checking algorithm that runs in polynomial time. It is further shown in [CDK93] that the CTL formulas they reduce to can be modified to formulas in a fragment slightly extending fair-CTL that can still be handled by the model checking algorithm of [EL87] for fair-CTL. Thus overall this gives a polynomial time algorithm for inclusion of DRAs and DMAs. Since our learning algorithm relies on the sample including shortlex examples and the algorithm above does not guarantee shortlex counterexamples, we give in Section 13 and Section 14 polynomial time automata-theoretic algorithms for testing inclusion for DRAs and DMAs that provide shortlex counterexamples, which are novel results.
@@ 12.2. An inclusion algorithm for DPAs
The inclusion problem for DPAs $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ over the same alphabet can be solved by looping over all odd $k_{1}$ in the range of $\kappa_{1}$ and all even $k_{2}$ in the range of $\kappa_{2}$, calling the $\mathbf{Colors}$ algorithm with inputs $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, $k_{1}$, and $k_{2}$. If the $\mathbf{Colors}$ algorithm returns any witness $u(v)^{\omega}$, then $u(v)^{\omega}\in{\llbracket}\mathcal{P}_{1}{\rrbracket}\setminus{\llbracket}\mathcal{P}_{2}{\rrbracket}$, and $u(v)^{\omega}$ is returned as a witness of non-inclusion. Otherwise, by Thm. 12.1, there is no $\omega$-word $w$ accepted by $\mathcal{P}_{1}$ and not accepted by $\mathcal{P}_{2}$, and the answer “yes” is returned for the inclusion problem. Note that for $i=1,2$, the range of $\kappa_{i}$ has at most $|Q_{i}|$ distinct elements. Thus we have the following.
@@
From Claims 1 and 4, we have the following.

### v4

@@ 8.4. The learning algorithm $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IPA}}}$
Let $C$ be any node of $\mathcal{F}^{*}(\mathcal{P})$, and let $D$ be a child of $C$ in $\mathcal{F}^{*}(\mathcal{P})$. Then $D$ is an SCC, $D\subseteq C$, the parity of $D$ is opposite to the parity of $C$, and $D$ is maximal in the subset ordering with these properties, by property (3) of Theorem 8.5. In the sample $T_{Acc}^{\mbox{\scriptsize{IPA}}}$ there is an example $(u(v)^{\omega},l)$ with $D={\textsl{inf}}(u(v)^{\omega})$, so $D$ is an element of $Z$, and will be made a child of $C$ in $\mathcal{F}^{\prime}$ because $D\subseteq C$, the label $l$ is the opposite of the label of $C$, and $D$ is maximal in $Z$ with these properties. Conversely, if $D$ is made a child of $C$ in $\mathcal{F}^{\prime}$, then $D\subseteq C$, the label of $D$ is opposite to the label of $C$ (that is, they are of opposite $\kappa$-parity), and $D$ is maximal in $Z$ with these properties. This implies $D$ is a child of $C$ in $\mathcal{F}^{*}(\mathcal{P})$, by property (3) of Theorem 8.5.
By induction, $\mathcal{F}^{\prime}$ is equal to $\mathcal{F}^{*}(\mathcal{P})$, and therefore $\kappa^{\prime}$ is equal to the canonical coloring $\kappa^{*}$. Then the IPA $(\mathcal{M},\kappa^{\prime})$ is equivalent to $\mathcal{P}$, by Theorem 8.6. Because $T$ is consistent with $\mathcal{P}$, the IPA $(\mathcal{M},\kappa^{\prime})$ is returned by $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IPA}}}$. ∎
@@
Assume the sample $T$ is consistent with $\mathcal{P}$ and subsumes $T_{L}$. By Theorem 6.2, the automaton $\mathcal{M}$ is isomorphic to the automaton of $\mathcal{P}$. By Theorem 8.7, the acceptor $(\mathcal{M},\kappa^{\prime})$ is equivalent to $\mathcal{P}$. Because it is consistent with $T$, the IPA $(\mathcal{M},\kappa^{\prime})$ is the acceptor returned by $\mathbf{L}^{\mbox{\scriptsize{IPA}}}$. ∎
9. The sample $T_{Acc}$ and learning algorithm for IRA
@@ 9.2. An ordering on sets of states
Let $\mathcal{M}=\langle\Sigma,Q,q_{\iota},\delta\rangle$ be a deterministic complete automaton in which every state is reachable from the initial state $q_{\iota}$. Recall from Section 6.2 that $\textit{access}(q)$ is the shortlex least string $s\in\Sigma^{*}$ such that $\delta(q_{\iota},s)=q$.
@@ 9.3. The learning algorithm $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IRA}}}$
Otherwise, $z_{i}$ causes the update to $\beta_{k}$, and among the pairs in $S_{k}$ that are added to $\beta_{k}$ there is at least one that $z_{i}$ satisfies. To see this, note that $z_{i}$ must satisfy some pair in $\alpha$, say $(q,B_{j})$. Thus, $q\in{\textsl{inf}}(z_{i})$ and $B_{j}\cap{\textsl{inf}}(z_{i})=\emptyset$. The pair $(q,B)$ where $B=Q\setminus{\textsl{inf}}(z_{i})$ has $B_{j}\subseteq B$. Thus any $\omega$-word that satisfies $(q,B)$ will also satisfy $(q,B_{j})$. Because $T$ is consistent with $\mathcal{R}$, there can be no negative example in $T$ satisfying $(q,B)$, so pair $(q,B)$ is part of $S_{k}$ and is added to $\beta_{k}$. The word $z_{i}$ (and the word $w$) satisfies $(q,B)$ and is therefore accepted by $(\mathcal{M},\beta_{k+1})$ and every subsequent hypothesis, including $(\mathcal{M},\beta)$. ∎
@@ 9.5. Correctness of $\mathbf{L}_{Acc}^{\mbox{\scriptsize{IRA}}}$
Now assume that the sample $T$ is consistent with $\mathcal{R}$ and subsumes $T_{L}$. Then by Theorem 6.2, the automaton $\mathcal{M}$ is isomorphic to the automaton of $\mathcal{R}$. By Theorem 9.4, the acceptor returned by $\mathbf{T}^{\mbox{\scriptsize{IRA}}}$ is an IRA $(\mathcal{M},\beta)$ that is equivalent to $\mathcal{R}$, and this is the acceptor also returned by $\mathbf{L}^{\mbox{\scriptsize{IRA}}}$. ∎
@@ 10.1. Computing $T_{Aut}$
If we have a procedure to solve the inclusion problem, at most two calls to it will solve the equivalence problem. We describe polynomial time algorithms to solve the inclusion problem for DBAs, DCAs and DPAs in Section 11 and 12, for DRAs and DSAs in Section 13, and for DMAs in Section 14. Referring to those sections, we obtain polynomial time algorithms to solve the equivalence problem for DBAs, DCAs and DPAs from Theorem 11.1 and Theorem 12.2, for DRAs and DSAs from Theorem 13.2, and for DMAs from Theorem 14.6. Thus, we have the following.
@@
Given an acceptor $\mathcal{A}$ of type IBA, ICA, IPA, IRA, ISA, or IMA, the sample $T_{Aut}$ for the automaton portion of $\mathcal{A}$ can be computed in polynomial time.
@@ 10.2. Computing $T_{Acc}$
Given an IBA $\mathcal{B}=\langle\Sigma,Q,q_{\iota},\delta,F\rangle$ with automaton $\mathcal{M}$, the sample $T_{Acc}^{\mbox{\scriptsize{IBA}}}$ described in Section 7.2 is computed as follows. For each SCC $C\in\textit{maxSCCs}(Q\setminus F)$, let $u(v)^{\omega}=\mathbf{Witness}(C,\mathcal{M})$ and include the negative example $(u(v)^{\omega},0)$ in $T_{Acc}^{\mbox{\scriptsize{IBA}}}$.
@@
Given an IPA $\mathcal{P}$ with automaton $\mathcal{M}$, the computation of $T_{Acc}^{\mbox{\scriptsize{IPA}}}$ proceeds as described in Section 8.3. That is, the canonical forest $\mathcal{F}^{*}(\mathcal{P})$ is computed in polynomial time, and for each node $C$ in the forest, $u(v)^{\omega}=\mathbf{Witness}(C,\mathcal{M})$ is computed and $(u(v)^{\omega},l)$ is added to $T_{Acc}^{\mbox{\scriptsize{IPA}}}$, where $l$ is the label of node $C$ in the canonical forest. ∎
@@
Given an IRA $\mathcal{R}=(\mathcal{M},\alpha)$, the computation of $T_{Acc}$ proceeds as described in Section 9.4. At each stage of the computation, it is necessary to find an $\omega$-word $u(v)^{\omega}$ with the $\preceq$-largest ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega}$) that is accepted by $(\mathcal{M},\alpha)$ and rejected by $(\mathcal{M},\gamma_{k})$. Theorem 13.2 gives a polynomial time algorithm that not only tests the inclusion of two DRAs, but returns a witness $u(v)^{\omega}$ with the $\preceq$-largest ${{\textsl{inf}}}_{\mathcal{M}}(u(v)^{\omega})$ in the case of non-inclusion, because $\mathcal{M}$ is isomorphic to $\mathcal{M}\times\mathcal{M}$. ∎
@@
Given an IMA $\mathcal{A}=(\mathcal{M},{\mathcal{F}})$, the sample $T_{Acc}^{\mbox{\scriptsize{IMA}}}$ described in Section 7.1 is computed as follows. For each $F\in{\mathcal{F}}$ determine whether $F$ is a reachable SCC of $\mathcal{M}$, and if so, compute $u(v)^{\omega}=\mathbf{Witness}(F,\mathcal{M})$ and add $(u(v)^{\omega},1)$ to the sample $T_{Acc}^{\mbox{\scriptsize{IMA}}}$. ∎
@@
Let $\mathcal{A}$ be an IBA, IPA, IRA, or IMA accepting the $\omega$-language $L$. Then the characteristic sample $T_{L}$ for $\mathcal{A}$ can be computed in polynomial time in the size of $\mathcal{A}$.
@@
By Theorem 10.1, the sample $T_{Aut}$ can be computed in polynomial time in the size of $\mathcal{A}$, and by Proposition 10.2, 10.3, 10.4, or 10.5 the sample $T_{Acc}^{\mbox{\scriptsize{IBA}}}$, $T_{Acc}^{\mbox{\scriptsize{IPA}}}$, $T_{Acc}^{\mbox{\scriptsize{IRA}}}$, or $T_{Acc}^{\mbox{\scriptsize{IMA}}}$ can also be computed in polynomial time in the size of $\mathcal{A}$. ∎
Note that Theorem 10.6 does not imply that the class $\mathbb{IXA}$ for $\mathbb{X}\in\{\mathbb{B},\mathbb{P},\mathbb{R},\mathbb{M}\}$ is efficiently teachable, since this class also has representations by non-isomorphic automata.
@@ 11. Inclusion algorithms
We show that there are polynomial time algorithms for the inclusion problem for DBAs, DCAs, DPAs, DRAs, DSAs and DMAs. Recall that two calls to an inclusion algorithm suffice to solve the equivalence problem. By Claim 2.2 (4), the inclusion and equivalence problems for DCAs are efficiently reducible to those for DBAs, and vice versa. Also, by Claim 2.2 (1), the inclusion and equivalence problems for DBAs are efficiently reducible to those for DPAs. Thus it suffices to consider the inclusion problem for DPAs, DRAs and DMAs.
@@
For the inclusion problem for DBAs, DCAs and DPAs, Schewe [Sch10, Sch11] gives the following result.
@@
The inclusion problems for DBAs, DCAs and DPAs are in NL.
Because NL (nondeterministic logarithmic space) is contained in polynomial time, this implies the existence of polynomial time inclusion and equivalence algorithms for DBAs, DCAs and DPAs. For the sake of completeness, and to address the problem of returning a witness we include a proof sketch.
@@
For inclusion of DRAs and DMAs, [CDK93] provides a reduction to the problem of model checking a formula in the temporal logic CTL. While the complexity of model checking CTL formulas is in general PSPACE-hard, for the fair-CTL fragment, [EL87] provides a model checking algorithm that runs in polynomial time. It is further shown in [CDK93] that the CTL formulas they reduce to can be modified to formulas in a fragment slightly extending fair-CTL that can still be handled by the model checking algorithm of [EL87] for fair-CTL. Thus overall this gives a polynomial time algorithm for inclusion of DRAs and DMAs. Since our learning algorithm relies on the sample including shortlex examples and the algorithm above does not guarantee shortlex counterexamples, we give in Section 13 and Section 14 polynomial time automata-theoretic algorithms for testing inclusion for DRAs and DMAs that provide shortlex counterexamples, which are novel results.
@@ 12.2. An inclusion algorithm for DPAs
The inclusion problem for DPAs $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ over the same alphabet can be solved by looping over all odd $k_{1}$ in the range of $\kappa_{1}$ and all even $k_{2}$ in the range of $\kappa_{2}$, calling the $\mathbf{Colors}$ algorithm with inputs $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, $k_{1}$, and $k_{2}$. If the $\mathbf{Colors}$ algorithm returns any witness $u(v)^{\omega}$, then $u(v)^{\omega}\in{\llbracket}\mathcal{P}_{1}{\rrbracket}\setminus{\llbracket}\mathcal{P}_{2}{\rrbracket}$, and $u(v)^{\omega}$ is returned as a witness of non-inclusion. Otherwise, by Theorem 12.1, there is no $\omega$-word $w$ accepted by $\mathcal{P}_{1}$ and not accepted by $\mathcal{P}_{2}$, and the answer “yes” is returned for the inclusion problem. Note that for $i=1,2$, the range of $\kappa_{i}$ has at most $|Q_{i}|$ distinct elements. Thus we have the following.
@@
From Claim 2.2 (1 and 4), we have the following.